OpenStudy (anonymous):

Choose the inequality that illustrates this problem. After serving 45 pounds of poultry, a caterer had more than 10 pounds left. How many pounds of poultry did the caterer have originally? Let p = pounds of poultry the caterer had originally @iGreen @jcpd910

OpenStudy (anonymous):

Okay, this is our equation: p - 45 = 10

OpenStudy (igreen):

Yep, now can you solve that? @temia12345

OpenStudy (anonymous):

This can be solved easily, but to prove it we must isolate p. p - 45 + 45 = 10 + 45

OpenStudy (anonymous):

equal to or less den

OpenStudy (anonymous):

or greater